OpenStudy (anonymous):

x-int: -3 y-int: 1 What is the equation of the line? I have the answer but i want to double check. I got y=1/3x+2

OpenStudy (anonymous):



OpenStudy (anonymous):

your eq. is wrong..

OpenStudy (anonymous):

slope intercept form i think you have it wrong

OpenStudy (anonymous):

y2-y1/x2-x1 then y-y1=m(x-x1)

OpenStudy (anonymous):

y=mx+c m=-(y/x) =-(1/-3) =1/3 at y=1,x=-3 point 1=1/3(-3)+c c=2 y=1/3+2 right answer

OpenStudy (anonymous):

y=1/3x+1

OpenStudy (anonymous):

why is it +1?

OpenStudy (anonymous):

m=1/3

OpenStudy (anonymous):

ok, when you put x=0 in your eq. you should get y=1 but you get 2

OpenStudy (anonymous):

1/3 x 3 = 3/3 = 1 then bring the other 1 then add and comes out to 2

OpenStudy (anonymous):

hmmmm